a man who can ride a bicycle in 50 miles in 2 hours and 25 minutes has what average speed in mph?
Answer by nerdybill(7384)

a man who can ride a bicycle in 50 miles in 2 hours and 25 minutes has what average speed in mph?
Thats 50 miles per 2hours and 25 minutes
.
First, convert
2 hours and 25 minutes
to "hours:
2 hours and 25 minutes = 2 + 25/60 = 2.4167 hours
.
50/2.4167 = 20.69 mph